# Unlocking General Long Chain-of-Thought Reasoning Capabilities of Large Language Models via Representation Engineering

### v1

@@ 1 Introduction
Recently, slow-thinking reasoning models, such as OpenAI’s o1 series of models OpenAI (2024) and DeepSeek-R1 Guo et al. (2025), have significantly advanced the capabilities of large language models (LLMs). As a typical approach, these reasoning models leverage long chain-of-thoughts (long CoTs), encompassing planning, validation, and backtracking strategies, to solve complex reasoning tasks Team et al. (2025); Yang et al. (2024); Pang et al. (2025). Most existing work focuses on eliciting long CoTs on tasks that are easy to verify, such as mathematics Yeo et al. (2025) and coding Xu et al. (2025). They find that the capability of long CoT reasoning can be efficiently elicited with only thousands of training examples Ye et al. (2025). Furthermore, some recent work finds that this capability can easily transfer to other tasks, even without any task-specific examples Du et al. (2025). These interesting phenomena raise a question: Is long CoT reasoning a general capability encoded in LLMs?
@@
In this work, we take the first step towards unraveling the mystery from the perspective of representation engineering Zou et al. (2023). As a transparent and interpretable method, representation engineering treats representation as the fundamental unit of analysis to understand and control high-level capabilities of LLMs, such as instruction following Stolfo et al. (2025), personality Cao et al. (2024), and hallucination Li et al. (2023); Arditi et al. (2024). Specifically, representations are extracted from the encodings of LLMs for data that reflect specific capabilities. These representations can then be used for analysis and control of model behaviors.
@@ 2 Related Work
Large Language Model Reasoning. Recently, improving the reasoning capabilities of LLMs has become a critical challenge. Prior approaches, such as test-time search Zhang et al. (2024); Guan et al. (2025), distillation Yu et al. (2024); Min et al. (2024), and reinforcement learning Guo et al. (2025); Team et al. (2025), enable LLMs to engage in deliberate thinking. Despite their remarkable success, the underlying mechanisms of LLM reasoning remain unclear. Some studies Christ et al. (2024); Rai and Yao (2024) analyze this by localizing specific neurons, but they only focus on isolated neuron connections, neglecting the cooperative activity of multiple neurons. Other work Hu et al. (2024); Højer et al. (2025) addresses this via representation engineering to better control neuron collaboration. However, these studies are often limited to short-form CoT, struggling to fully unlock the reasoning potential of LLMs. In this paper, we focus on exploring the mechanism of long CoT reasoning through representation engineering.
Representation Engineering. Representation engineering treats internal representations as the fundamental unit, focusing on analyzing and manipulating them within neural networks. As a well-established technique, it has been applied in various areas such as personality modeling Cao et al. (2024), instruction following Stolfo et al. (2025), hallucination alleviation Li et al. (2023); Arditi et al. (2024), and safety improvement Liu et al. (2024). While prior work focuses on simple concepts like sentiment Hollinsworth et al. (2024) and style von Rütte et al. (2024); Scalena et al. (2024), our work aims to address the more complex challenge: understanding and unlocking general long CoT reasoning capabilities of LLMs.
@@ 3.2 Analysis of Long CoT Representations
Analysis of General Representations. To analyze the characteristics of vanilla and long CoTs, we visualize their representations to compare their distributions. Specifically, we employ a dimensionality reduction approach (i.e., t-SNE van der Maaten and Hinton (2008)) to map representations obtained from the above part onto a 2D plane. As illustrated in Figure 2(a) (more figures in Appendix D), the representations of various long CoTs are concentrated in a specific area of the whole space. In addition, their distribution areas are clearly distinct from those of vanilla CoTs. Taken together, the two pieces of evidence suggest that LLMs do encode long CoT reasoning as a separate general capability in their parameter spaces. Moreover, we find that the separation between these two types of CoTs is the most pronounced in the middle layers of the model, while less clear in the early and final layers. This phenomenon may be attributed to the fact that middle layers integrate information from early layers and are more informative Skean et al. (2025), playing a critical role in capturing high-level concepts (e.g., CoT reasoning) (See Appendix A).
Analysis of Domain-Specific Representations. In this part, we further examine the characteristics of vanilla and long CoTs in specific domains. Specifically, we collect representations in mathematical and other domains (i.e., physics, chemistry, and biology) and visualize them following the previous part. As shown in Figure 2(b) (more figures in Appendix E), different domains share similar contrastive representations between long and vanilla CoTs, which further demonstrates the transferability of long CoT reasoning. In addition, the representations of mathematical domains are relatively concentrated, while those of other domains (e.g., physics) are more dispersed. This may be due to the fact that mathematical problems focus on logical reasoning patterns, while problems in other domains also require domain-specific information. That is, domain-specific CoT data plays an important role for the elicitation of long CoT reasoning within these domains.
@@ 5.1 Experimental Setup
Baselines. To facilitate a systematic comparison, we select several representative methods, including prompting-based approaches (i.e., Zero-shot CoT, Few-shot CoT, and BoostStep Zhang et al. (2025)), neuron activation method (i.e., MathNeuro Christ et al. (2024)), representation engineering method (i.e., RoT Hu et al. (2024)), and supervised fine-tuning method. Detailed descriptions of these baselines are provided in Appendix F.
Implementation Details. In our experiments, we use two representative open-source LLMs: Qwen2.5-7B-Instruct Yang et al. (2024) and Llama3.1-8B-Instruct Dubey et al. (2024). For the contrastive reasoning pattern representation, we set the injection strength $\lambda_{p}$ as 0.1. For the question-aware domain-specific representation, we set the number of retrieved representations $k$ as 8 and the injection strength $\lambda_{d}$ as 0.1. Both representations are injected into the intermediate layer of LLMs.
@@ 5.2 Experimental Results
The experimental results are presented in Table 1. As we can see, the few-shot CoT method (i.e., few-shot CoT and BoostStep) performs poorly, even worse than the zero-shot CoT. The main reason is that LLMs lose their ability to learn from demonstrations after supervised fine-tuning Wei et al. (2023). Adding examples at the discrete token level can increase input length, which distracts the model from the current problem and disrupts the reasoning process. In contrast, MathNeuro can improve performance by identifying and scaling relevant neurons, but they focus solely on specific neuron connections, neglecting the cooperative activity of multiple neurons. To address this limitation, RoT introduces contrastive representations of CoT and non-CoT prompts, enabling fine-grained control over the reasoning process. However, this approach is insufficient to guide the model toward deliberate thinking, resulting in limited performance improvements.
@@ 5.3.2 The Efficiency of GLoRE
In this part, We discuss the efficiency of GLoRE, as shown in Table 3. First, the few-shot CoT method incorporates additional demonstrations in the input, which leads to the increased inference time. This is because the time complexity of the Transformer scales quadratically with the length of the input sequence. Additionally, BoostStep decomposes the reasoning process into multiple substeps and guides the model with relevant examples at each step. This requires the model to perform multiple reasoning iterations for a single problem, further increasing the computational overhead. In contrast, GLoRE maintains the same time complexity as zero-shot CoT, which is significantly lower than few-shot CoT and BoostStep. We extract representation from the model’s latent space and inject them during inference, without reducing the reasoning efficiency. This demonstrates that GLoRE can significantly enhance reasoning capabilities through representation engineering while preserving computational efficiency.
@@ 7 Limitations
The matrix-based entropy metrics for vanilla and long CoTs representations across different layers in Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct are illustrated in Figure 8. We observe that the matrix-based entropy of long CoT is consistently higher than that of vanilla CoT, indicating that long CoT contains more diverse and less redundant features within the latent space. Additionally, we find that the entropy in the middle layers of the model is higher than in the final layer in both CoTs. This suggests that the middle layers are better at extracting diverse and complex features, exhibiting powerful capabilities in reasoning tasks El-Nouby et al. (2024); Fan et al. (2024).
Appendix B More experiments
B.1 The Effect of the Representation Memory Size
(a) MATHOAI
(b) GPQA
In this part, we investigate the impact of scaling the representation memory size on GLoRE. The results are illustrated in 9. As we can see, the performance consistently improves as the number of examples in the representation memory increases. This improvement can be attributed to two main reasons. On one hand, when computing the average representation of the contrastive reasoning pattern representation, using a larger number of demonstrations helps better isolate problem-specific information, resulting in more precise high-level long CoT pattern representations. On the other hand, when extracting and injecting domain-specific features, the increased representation memory provides the model with access to more relevant question-aware domain-specific representations, enabling fine-grained refinement of the reasoning process in specific domains.
B.2 The Effect of the Representation Injection Position
CRR DR MATHOAI GPQA
First First 74.40 35.86
First Last 76.20 36.36
Last First 75.80 35.86
Last Last 75.00 34.85
Table 4: Performance comparison across different injection positions on MATHOAI and GPQA on Qwen2.5-7B-Instruct. “CRR” and “DR” denote the Contrastive Reasoning Pattern Representation and Question-Aware Domain-Specific Representation, respectively.
In this part, we investigate the effect of the representation injection position. As illustrated in Table 4, we find that they have minimal effect on performance with the best results achieved when the contrastive reasoning pattern representation and the domain-specific representation are injected into the model’s first and last tokens, respectively.
B.3 The Impact of Layer Selection
Appendix C Statistics of vanilla and long CoT Examples
@@
Appendix D Detailed Visualization of Short-form and Long-form Thought Representations
@@
Appendix E Detailed Visualization of Representations across Different Domains
@@
Appendix F Detailed Description of Baselines.

### v2

@@ 1 Introduction
Recently, slow-thinking reasoning models, such as OpenAI’s o1 series of models OpenAI (2024) and DeepSeek-R1 Guo et al. (2025), have significantly advanced the capabilities of large language models (LLMs) Zhao et al. (2023). As a typical approach, these reasoning models leverage long chain-of-thoughts (long CoTs), encompassing planning, validation, and backtracking strategies, to solve complex reasoning tasks Yang et al. (2024); Team et al. (2025); Pang et al. (2025). Most existing work focuses on eliciting long CoTs on tasks that are easy to verify, such as mathematics Cheng et al. (2024a); Yeo et al. (2025) and coding Xu et al. (2025). They find that the capability of long CoT reasoning can be efficiently elicited with only thousands of training examples Ye et al. (2025). Furthermore, some recent work finds that this capability can easily transfer to other tasks, even without any task-specific examples Du et al. (2025). These interesting phenomena raise a question: Is long CoT reasoning a general capability encoded in LLMs?
@@
In this work, we take the first step towards unraveling the mystery from the perspective of representation engineering Zou et al. (2023). As a transparent and interpretable method, representation engineering treats representation as the fundamental unit of analysis to understand and control high-level capabilities of LLMs, such as instruction following Stolfo et al. (2025), personality Cao et al. (2024), and hallucination Li et al. (2023a); Arditi et al. (2024). Specifically, representations are extracted from the encodings of LLMs for data that reflect specific capabilities Dong et al. (2024a). These representations can then be used for analysis and control of model behaviors.
@@ 2 Related Work
Large Language Model Reasoning. Recently, improving the reasoning capabilities of LLMs has become a critical challenge. Prior approaches, such as test-time search Zhang et al. (2024); Tang et al. (2024b); Guan et al. (2025); Cheng et al. (2025), distillation Yu et al. (2024); Min et al. (2024), and reinforcement learning Guo et al. (2025), enable LLMs to engage in deliberate thinking Tang et al. (2024a); Cheng et al. (2024a); Wang et al. (2025a). Despite their remarkable success, the underlying mechanisms of LLM reasoning remain unclear. Some studies Christ et al. (2024); Rai and Yao (2024) analyze this by localizing specific neurons, but they only focus on isolated neuron connections, neglecting the cooperative activity of multiple neurons. Other work Hu et al. (2024); Højer et al. (2025) addresses this via representation engineering to better control neuron collaboration. However, these studies are often limited to short-form CoT, struggling to fully unlock the reasoning potential of LLMs. In this paper, we focus on exploring the mechanism of long CoT reasoning through representation engineering.
Representation Engineering. Representation engineering Zou et al. (2023) treats internal representations as the fundamental unit, focusing on analyzing and manipulating them within neural networks. As a well-established technique, it has been applied in various areas such as personality modeling Cao et al. (2024), instruction following Stolfo et al. (2025), hallucination alleviation Li et al. (2023a); Arditi et al. (2024); Li et al. (2024), and safety improvement Liu et al. (2024). While prior work focuses on simple concepts like sentiment Hollinsworth et al. (2024) and style von Rütte et al. (2024); Scalena et al. (2024), our work aims to address the more complex challenge: understanding and unlocking general long CoT reasoning capabilities of LLMs.
@@ 3.2 Analysis of Long CoT Representations
Analysis of General Representations. To analyze the characteristics of vanilla and long CoTs, we visualize their representations to compare their distributions. Specifically, we employ a dimensionality reduction approach (i.e., t-SNE van der Maaten and Hinton (2008)) to map representations obtained from the above part onto a 2D plane. As illustrated in Figure 2(a) (more figures in Appendix C), the representations of various long CoTs are concentrated in a specific area of the whole space. In addition, their distribution areas are clearly distinct from those of vanilla CoTs. Taken together, the two pieces of evidence suggest that LLMs do encode long CoT reasoning as a separate general capability in their parameter spaces. Moreover, we find that the separation between these two types of CoTs is the most pronounced in the middle layers of the model, while less clear in the early and final layers. This phenomenon may be attributed to the fact that middle layers integrate information from early layers and are more informative Skean et al. (2025), playing a critical role in capturing high-level concepts (e.g., CoT reasoning) (See Appendix A).
Analysis of Domain-Specific Representations. In this part, we further examine the characteristics of vanilla and long CoTs in specific domains. Specifically, we collect representations in mathematical and other domains (i.e., physics, chemistry, and biology) and visualize them following the previous part. As shown in Figure 2(b) (more figures in Appendix D), different domains share similar contrastive representations between long and vanilla CoTs, which further demonstrates the transferability of long CoT reasoning. In addition, the representations of mathematical domains are relatively concentrated, while those of other domains (e.g., physics) are more dispersed. This may be due to the fact that mathematical problems focus on logical reasoning patterns, while problems in other domains also require domain-specific information. That is, domain-specific CoT data plays an important role for the elicitation of long CoT reasoning within these domains Dong et al. (2025b).
@@ 5.1 Experimental Setup
Baselines. To facilitate a systematic comparison, we select several representative methods, including prompting-based approaches (i.e., Zero-shot CoT, Few-shot CoT, and BoostStep Zhang et al. (2025)), neuron activation method (i.e., MathNeuro Christ et al. (2024)), representation engineering method (i.e., RoT Hu et al. (2024)), and supervised fine-tuning method. Detailed descriptions of these baselines are provided in Appendix E.
Implementation Details. In our experiments, we use two representative open-source LLMs: Qwen2.5-7B-Instruct Yang et al. (2024) and Llama3.1-8B-Instruct Dubey et al. (2024). For the contrastive reasoning pattern representation, we set the injection strength $\lambda_{p}$ as 0.1. For the question-aware domain-specific representation, we set the number of retrieved representations $k$ as 8 and the injection strength $\lambda_{d}$ as 0.1. Both representations are injected into the intermediate layer of LLMs. Following the existing work Li et al. (2023b); Wang et al. (2024); Cheng et al. (2024b), we use the greedy decoding strategy for inference.
@@ 5.2 Experimental Results
The experimental results are presented in Table 1. As we can see, the few-shot CoT method (i.e., few-shot CoT and BoostStep) performs poorly, even worse than the zero-shot CoT. The main reason is that LLMs lose their ability to learn from demonstrations after supervised fine-tuning Wei et al. (2023). Adding examples at the discrete token level can increase input length, which distracts the model from the current problem and disrupts the inference process Dong et al. (2024b); Peng et al. (2025). In contrast, MathNeuro can improve performance by identifying and scaling relevant neurons, but it focuses solely on specific neuron connections, neglecting the cooperative activity of multiple neurons. To address this limitation, RoT introduces contrastive representations of CoT and non-CoT prompts, enabling fine-grained control over the reasoning process. However, this approach is insufficient to guide the model toward deliberate thinking, resulting in limited performance improvements.
@@ 5.3.2 The Efficiency of GLoRE
In this part, We discuss the efficiency of GLoRE, as shown in Table 3. First, the few-shot CoT method incorporates additional demonstrations in the input, which leads to the increased inference time. This is because the time complexity of the Transformer scales quadratically with the length of the input sequence Dong et al. (2025a); Zhan et al. (2024); Liu et al. (2025). Additionally, BoostStep decomposes the reasoning process into multiple substeps and guides the model with relevant examples at each step. This requires the model to perform multiple reasoning iterations for a single problem, further increasing the computational overhead. In contrast, GLoRE maintains the same time complexity as zero-shot CoT, which is significantly lower than few-shot CoT and BoostStep. We extract representation from the model’s latent space and inject them during inference, without reducing the reasoning efficiency. This demonstrates that GLoRE can significantly enhance reasoning capabilities through representation engineering while preserving computational efficiency.
@@ 5.3.4 Experiments on Larger Models
5.3.5 The Effect of the Representation Memory Size
(a) MATHOAI
(b) GPQA
In this part, we investigate the impact of scaling the representation memory size on GLoRE. The results are illustrated in 8. As we can see, the performance consistently improves as the number of examples in the representation memory increases. This improvement can be attributed to two main reasons. On one hand, when computing the average representation of the contrastive reasoning pattern representation, using a larger number of demonstrations helps better isolate problem-specific information, resulting in more precise high-level long CoT pattern representations. On the other hand, when extracting and injecting domain-specific features, the increased representation memory provides the model with access to more relevant question-aware domain-specific representations, enabling fine-grained refinement of the reasoning process in specific domains.
5.3.6 The Impact of Layer Selection
@@ 7 Limitations
Acknowledgements
This work was partially supported by National Natural Science Foundation of China under Grant No. 92470205 and 62222215, Beijing Municipal Science and Technology Project under Grant No. Z231100010323009, Beijing Natural Science Foundation under Grant No. L233008, and Ant Group. Xin Zhao is the corresponding author.
@@ Acknowledgements
The matrix-based entropy metrics for vanilla and long CoTs representations across different layers in Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct are illustrated in Figure 10. We observe that the matrix-based entropy of long CoT is consistently higher than that of vanilla CoT, indicating that long CoT contains more diverse and less redundant features within the latent space. Additionally, we find that the entropy in the middle layers of the model is higher than in the final layer in both CoTs. This suggests that the middle layers are better at extracting diverse and complex features Wang et al. (2025b), exhibiting powerful capabilities in reasoning tasks El-Nouby et al. (2024); Fan et al. (2024).
Appendix B Statistics of vanilla and long CoT Examples
@@
Appendix C Detailed Visualization of Short-form and Long-form Thought Representations
@@
Appendix D Detailed Visualization of Representations across Different Domains
@@
Appendix E Detailed Description of Baselines.